What's the answer to 32n-9=(4n+3)?
Answer by algebrahouse.com(1659)

32n - 9 = 4n + 3
28n - 9 = 3 {subtracted 4n from each side}
28n = 12 {added 9 to each side}
n = 12/28 {divided each side by 28}
n = 3/7 {reduced by a factor of 4}

For more help from me, visit: www.algebrahouse.com